OpenStudy (anonymous):

.02(6+-3)=.12(+-2)+.18 i tried to solve it butt i cant get the correct anser the correct anser is all real numbers

OpenStudy (anonymous):

what is the variable here?

OpenStudy (anonymous):

sorry let me re type it .02(6t-3)=.12(t-2)+.18

OpenStudy (anonymous):

you are not getting an answer because it is an identity. true for all t

OpenStudy (anonymous):

i.e. the thing on the right is always equal to the thing on the left

OpenStudy (anonymous):

but i dont know how to back it up with the work i tryed to solve it and i get crazy answere everytime i try to solve it

OpenStudy (anonymous):

wait there is nothing to solve the left hand side is exactly the same as the right had side

OpenStudy (anonymous):

like if i try to solve \[x+3=x+3\] it is the same thing on both sides so there is on one solution any number works

OpenStudy (anonymous):

the left hand side is \[.12t-.06\] when you multiply out

OpenStudy (anonymous):

the right hand side is \[.12t-.24+.18=.12t-.06\] so both sides are the same

OpenStudy (anonymous):

omg i was adding insted of multiplying this whole time just catched it cuz of you thanks so much !

OpenStudy (anonymous):

yw